40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 298 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  2180   Thu Nov 5 16:24:40 2009 JenneUpdateGeneralDrill Press is down for the count

The on/off switch for the drill press is broken.  Replacement parts should be here tomorrow. 

  2179   Thu Nov 5 12:34:26 2009 kiwamuUpdateComputerselog rebooted

I found elog got crashed. I rebooted the elog daemon just 10minutes before.

  2178   Thu Nov 5 05:07:22 2009 ranaUpdateLSCX Arm Cavity transfer Function

I would have guessed that you have to calibrate the detectors relative to each other before trying this. Its also going to be tricky if you use 2 different kinds of ADC for this (c.f. today's delay discussion in the group meeting).

I think Osamu used to look at fast transmission signals by making sure the PD at the end had a 50 Ohm output impedance and just drive the 40m long cable and terminate the receiving end with 50 Ohms. Then both PDs go into the SR785.

 

  2177   Wed Nov 4 23:17:51 2009 AlbertoUpdateLSCX Arm Cavity transfer Function

I measured the transfer function between MC_TRANS and TRX and I'm attaching the result.

ArmCavityTF02.png

That looks quite strange. Something's wrong. I'll repeat it tomorrow.

for the night I'm putting everything back. I'm also reconnecting the OMC_ISS_EXC and opening again the test switch on the ISS screen.

The RFAM monitor remains disable

  2176   Wed Nov 4 21:46:18 2009 AlbertoUpdateLSCArm Cavity Finesse Measurement

Quote:

Quote:

I'm going to work on the X arm to measure the arm cavity finesse.

The idea is to measure the cavity transfer function to estimate the frequency of its cavity pole. That should be a more accurate measurement than that based on the cavity decay time.

I'm starting now and I'm going to inject a swept sine excitation on the OMC_ISS_EXC input cable laying on the floor nearby the AP table (see pic).

DSC_0952.JPG

In orderf to do that I disconnected the cable from the OMC breakout box laying on the floor. I'm going to plug the cable back in as soon as I'm done.

 Since I need to measure the transfer function between TRX and MC_TRANS_DC I picked off the beam going to RFAM PD to send it to a PDA255 photodiode (cannibalized from the AbsL's PLL) which I installed on the PSL table.

I centerd the beam on the PD and I was able to see the injected signal.

I think I'm ready to measure the transfer function.

Except for the RFAM PD everything is as before.

I'm gonna go grab dinner and I should be back to keep working on that in about one hour.

 Back from dinner. Taking measurements.

  2175   Wed Nov 4 18:35:19 2009 AlbertoUpdateLSCArm Cavity Finesse Measurement

Quote:

I'm going to work on the X arm to measure the arm cavity finesse.

The idea is to measure the cavity transfer function to estimate the frequency of its cavity pole. That should be a more accurate measurement than that based on the cavity decay time.

I'm starting now and I'm going to inject a swept sine excitation on the OMC_ISS_EXC input cable laying on the floor nearby the AP table (see pic).

DSC_0952.JPG

In orderf to do that I disconnected the cable from the OMC breakout box laying on the floor. I'm going to plug the cable back in as soon as I'm done.

 Since I need to measure the transfer function between TRX and MC_TRANS_DC I picked off the beam going to RFAM PD to send it to a PDA255 photodiode (cannibalized from the AbsL's PLL) which I installed on the PSL table.

I centerd the beam on the PD and I was able to see the injected signal.

I think I'm ready to measure the transfer function.

Except for the RFAM PD everything is as before.

I'm gonna go grab dinner and I should be back to keep working on that in about one hour.

  2174   Wed Nov 4 16:49:32 2009 AlbertoUpdateLSCArm Cavity Finesse Measurement

I'm going to work on the X arm to measure the arm cavity finesse.

The idea is to measure the cavity transfer function to estimate the frequency of its cavity pole. That should be a more accurate measurement than that based on the cavity decay time.

I'm starting now and I'm going to inject a swept sine excitation on the OMC_ISS_EXC input cable laying on the floor nearby the AP table (see pic).

DSC_0952.JPG

In orderf to do that I disconnected the cable from the OMC breakout box laying on the floor. I'm going to plug the cable back in as soon as I'm done.

  2173   Tue Nov 3 12:47:01 2009 KojiConfigurationCDS1Y9 Rack configuration update

For the CDS upgrade preparation I put and moved those stuff at the rack 1Y9:

Placed 1Y9-12 ADC to DB44/37 Adapter LIGO D080397

Placed 1Y9-14 DAC to IDC Adapter LIGO D080303

Moved the ethernet switch from 1Y9-16 to 1Y9-24

Wiki has also been updated.

  2172   Tue Nov 3 03:45:04 2009 rob UpdateIOOfrequency noise problem

Quote:

I used the XARM as a reference to measure the frequency noise after the MC.  It's huge around 4kHz--hundreds of times larger than the frequency noise the MC servo is actually squashing.  This presents a real problem for our noise performance.

An elog search reveals that this noise has been present (although not calibrated till now) for years.  We're not sure what's causing it, but suspicion falls on the piezojena input PZTs. 

I didn't bother too much about it before because we previously had enough common mode servo oomph to squash it below other DARM noises, and I didn't worry too much about stuff at 4kHz..  Now that we have a weaker FSS and thus much weaker CM servo, we can't squash it, and the most interesting feature of our IFO is at 4kHz. 

I'll measure the actual voltage noise going to the PZTs.  I remember doing this before and concluding it was ok, but can't find an elog entry.  So this time maybe I'll  do it right.

 

This level of frequency noise has not changed, but we now have increased common mode servo gain and so it's not as huge of a deal, although we should still probably do something about it. 

 

Attached is a plot of the piezojena noise measurement, estimated into Hz, along with another measurement of frequency noise as described above. 

To get the piezojena voltage noise into Hz, I estimated the PZTs within have a flat 2 micron/V response (based on a rough knowledge of their geometry and assuming a 10 milliradian / 150V steering range).  This is the voltage noise with the PZTs operating in closed loop mode, which is how we normally run them.  This plot also ignores the transfer function of the Pomona box, as we are mainly looking at noise in the kHz band.  I think this plot shows that these PZTs are a good candidate for creating this frequency noise, especially near their mechanical resonances (the manual says the unloaded resonances are in the 3-4kHz range).   

I've been operating one DOF of the piezojenas in open loop mode for a couple of weeks now, and the feared drift has not been a problem at all.  If we plan to keep using these after the upgrade, we should definitely put some big resistors in series at the outputs and operate them in open loop mode.

Also attached is a plot of RF DARM noise, with a frequency noise spectrum.  That spectrum is a REFL 2I spectrum put into DARM units using a measured TF (driving MC_L and measuring REFL 2I and DARM_ERR), and then put into meters using the same DARM calibration as used for the DARM curve.

Attachment 1: noise.png
noise.png
Attachment 2: spectra.pdf
spectra.pdf
  2171   Mon Nov 2 21:09:15 2009 SanjitUpdateAdaptive FilteringMore work on saving coeffs on the OAF screen

 

I made some changes in the code (all commented in the installed and SVN version) to print the filter coefficients. I got crazy output. Sometimes memory bugs lead to such crazy behavior. So far I could not find any bugs, but will have to spend more time on it.

 

  2170   Mon Nov 2 15:27:08 2009 JenneUpdatePEMGur2 cables have been moved

The cables labeled "Gur2" which were connected to channels 2,3,4 of the PEM-ADCU have been moved to the PCIX ADC which is connected directly to the ASS machine.  This means that until I (a) put the cables back or (b) figure out how to route channels from the ASS ADC to the RFM, we won't be able to use these channels for environmental monitoring, nor will they be saved. 

The Gur2_X, Gur2_Y, Gur2_Z channels are now connected to the 2nd, 3rd and 4th ADC channels respectively, on the ASS ADC (the first channel / TP1 is ADC0_0, which is the 1pps signal.).  The sketchy thing about the setup is the connection between the cables and the new ADC board.  The PCIX card is connected to the ASS via a white ribbon cable, and the board is just sitting on top of the computer box.  The 1pps (which has been hooked up for a long time) goes into the board via clip-doodles.  The regular channels have a SCSI cable connector, so I used a SCSI cable to connect up the ADC tester board, and connected my seismometer inputs to this tester board via more clip doodles.  Clearly this is a sketchy solution, and not okay for more than a day or so.  But we'll see how it goes. 

I'm going to, on the SimuLink Diagram, change the input source of these channels, from the RFMN to the ADC.  Then we'll see if that fixes our timing problem, and magically makes everything relating to the OAF work, and subtract huge amounts of noise. 

  2169   Mon Nov 2 13:34:36 2009 kiwamuConfigurationPSLremoved multiply resonant EOM

I removed the multiply resonant EOM that has been set by a SURF student from PSL table.

I will use it for checking the resonant circuit.

  2168   Mon Nov 2 13:00:55 2009 KojiUpdateIOOPMC aligned, MC WFS aligned

The beam to PMC aligned. The beam to MC WFS cameras aligned.
PMC Trans increased from 2.73 to 2.75 (+1%).
MC Trans increased from 7.80 to 7.87 (+1%).

  2167   Mon Nov 2 10:56:09 2009 kiwamuUpdateLSCcron job works succesfully & no timing jitter

As I wrote on Oct.27th, the cron job works every Sunday.

I found it worked well on the last Sunday (Nov.1st).

And I can not find any timing jitter in the data, its delay still stay 3*Ts.

  2166   Sun Nov 1 17:58:44 2009 JenneUpdateGeneralUpdate on Video Switch

The current update on the Chameleon video switch is: no progress.

I connected the old laptop that Rob/Steve acquired via RS-232 serial to the back of the video switch.  I'm using P2, the same serial port that the C1AUX computer was connected to just in case there's something good about P2 vs. P1. 

I used HyperTerminal to (try to) talk to the switch.  Settings were:  COM1, bits per second = 9600, data bits = 8, parity = none, stop bits = 1, flow control = none.  I can successfully send/get back responses to the basic commands, I (inquiry as to the type of equipment), and H (help - spits out the list of acceptable commands).  But when I try to do an actual command to do some video switching, everything hangs.  The front panel's rolling display (which just echos the company name) stops, then starts up again after ~20sec.  The hyperterminal display doesn't change.  I get neither the "DONE" answerback, which would indicate that the command executed successfully, nor do I get the "ERROR" answerback, which would indicate that something is wrong.  It just hangs.  If I disconnect, and restart the connection, and instead of trying a real command, but instead just send 'blahblahblah', then it will answerback 'ERROR' the first time, and then if I try to send another garbage message, everything hangs again.  So, I can sort of talk to the video switch, but I can't make it do anything yet.

I'm leaving the laptop connected instead of C1AUX, since the video EPICS screen doesn't work anyway for now.  If you want to start up the connection, either input the settings quoted above, or open "40m Video", which should have these connection settings saved in HyperTerminal.

  2165   Fri Oct 30 10:52:56 2009 JenneUpdatePSLHEPAs

Zach found the HEPA switch on the PSL table OFF.  He turned them on.

  2164   Fri Oct 30 09:24:45 2009 steveHowToMOPAhow to squeeze more out of little

Quote:

Here is the plots for the powers. MC TRANS is still rising.

What I noticed was that C1:PSL-FSS_PCDRIVE nolonger hit the yellow alert.
The mean reduced from 0.4 to 0.3. This is good, at least for now.

 Koji did a nice job increasing light power with some joggling.

Attachment 1: 44to34.jpg
44to34.jpg
  2163   Fri Oct 30 04:41:37 2009 robUpdateLockingworking again

I never actually figured out exactly what was wrong in entry 2162, but I managed to circumvent by changing the time sequence of events in the up script, moving the big gain increases in the common mode servo to the end of the script.  So the IFO can be locked again.

  2162   Thu Oct 29 21:51:07 2009 robUpdateLockingbad

Quote:

Quote:

Lock acquisition has gone bad tonight. 

The initial stage works fine, up through handing off control of CARM to MCL.  However, when increasing the AO path (analog gain), there are large DC shifts in the C1:IOO-MC_F signal.  Eventually this causes the pockels cell in the FSS loop to saturate, and lock is lost. 

 This problem has disappeared.  I don't know what it was. 

The first plot shows one of the symptoms.  The second plot is a similar section taken from a more normal acquisition sequence the night before.

All is not perfect, however, as now the handoff to RF CARM is not working.

 

The problem has returned.  I still don't know what it is, but it's making me angry. 

Attachment 1: itsback.png
itsback.png
  2161   Thu Oct 29 20:21:14 2009 KojiUpdatePSLNPRO LTMP lowered 9.5deg

Here is the plots for the powers. MC TRANS is still rising.

What I noticed was that C1:PSL-FSS_PCDRIVE nolonger hit the yellow alert.
The mean reduced from 0.4 to 0.3. This is good, at least for now.

Attachment 1: PSL_MC.png
PSL_MC.png
  2160   Thu Oct 29 18:25:33 2009 SanjitUpdateAdaptive FilteringMore work on saving coeffs on the OAF screen

Quote:

[Sanjit,Jenne]

Sanjit has been working today on trying to get the OAF coefficients to save properly.  Alex got us most of the way, but right now it's looking like the filter that is being saved is totally constant (all the values are the same).  We're poking around trying to figure out why this is. 

Also, we're starting again (as we should have been for the last week or so since Alex came in to help us) to check in the TOP_XFCODE whenever we make changes to it, and when we recompile the front end code. 

 

We are manually restarting assepics, but the terminal logs us out after sometime and ass may crash. I set autologout=0 in the terminal for the time being. Once the testing process is over, assepics will start automatically when the computer is turned on, so we wont have to worry about this.

(if ass crashes tonight, it is not unexpected!)

 

  2159   Thu Oct 29 18:04:02 2009 JenneUpdateAdaptive FilteringMore work on saving coeffs on the OAF screen

[Sanjit,Jenne]

Sanjit has been working today on trying to get the OAF coefficients to save properly.  Alex got us most of the way, but right now it's looking like the filter that is being saved is totally constant (all the values are the same).  We're poking around trying to figure out why this is. 

Also, we're starting again (as we should have been for the last week or so since Alex came in to help us) to check in the TOP_XFCODE whenever we make changes to it, and when we recompile the front end code. 

  2158   Thu Oct 29 13:48:32 2009 KojiUpdatePSLNPRO LTMP lowered 9.5deg

13:00 Found MC TRANS less than 7.
13:50 Go into the PSL table.
14:20 Work done. Now I am running SLOWscan script.
15:10 SLOWscan finished. It was not satisfactory. I go into the table again.
15:15 Running SLOWscan again.
16:00 SLOWscan done. Lock PMC. Adjust NPRO current so as to maximize PMC TRANS.
16:10 Lock RC, PMC, MZ, MC. Align PMC / MZ on the table. Align MC WFS beams on the QPDs.
16:30 Work done.

New FSS-SLOWDC nominal is -4.0

Now MC TRANS is 7.9. This is +12% increase. ENJOY!
HEPA is on at 90%. Light is off.

---------

NPRO TEMP trimmer adjustment
o PSL NPRO TEMP trimmer at the back of the laser head was turned 6.5 times in CW.
o It reduced NPRO crystal temp by 9.5deg. (43.5deg -> 34.0deg for FSS_SLOWDC -5.5)

To revert the previous setting, refer to the former measurement
c.f. http://nodus.ligo.caltech.edu:8080/40m/2008

NPRO Thermal scan
o 2 scans are performed.
o I selected the colder side of the second scan. i.e. SLOWDC=-4.0

NPRO Current adjustment
o Tweaked C1:PSL-126MOPA_126CURADJ while looking at PMC TRANS.
o CURADJ was changed from -2.25 to -1.9. This corresponds to change of C1:PSL-126MOPA_CURMON from 2.503A to 2.547A.

Attachment 1: 091028_PSL.png
091028_PSL.png
  2157   Wed Oct 28 17:20:21 2009 ranaSummaryCOCETM HR reflectivity plot

This is a plot of the R and T of the existing ETM's HR coating. I have only used 1/4 wave layers (in addition to the standard 1/2 wave SiO2 cap on the top) to get the required T.

The spec is a T = 15 ppm +/- 5 ppm. The calculation gives 8 ppm which is close enough. The calculated reflectivity for 532 nm is 3%. If the ITM reflectivity is similar, the signal for the 532 nm locking of the arm would look like a Michelson using the existing optics.

etm_40_1998.png

  2156   Wed Oct 28 14:39:10 2009 steveUpdatePEMafter the tour of the 40m
Illuminators and PSL lights turned off.
HEPA filter speed increased from 20 to 100%
  2155   Wed Oct 28 09:12:18 2009 steveUpdatePSLPMC power on the rise?

The PMC power is seems to be on the rise, ( MOPA_AMPMON is dropping ?) but I do not think it is real. We have Santa Anna wind condition, when the relative humidity drops and  ......

There is an other funky think. The room temp became rock solid. The PSL HEPAs running at 20% and IFO-room ACs are also in normal operational mode.

Attachment 1: pmcprising.jpg
pmcprising.jpg
  2154   Wed Oct 28 05:02:28 2009 robUpdateLockingback

LockAcq is back on track, with the full script working well.  Measurements in progress.

  2153   Tue Oct 27 19:37:03 2009 kiwamuUpdateLSCcron job to diagnose LSC-timing

I set a cron job on allegra.martian to run the diagnostic script every weekend.

I think this routine can be helpful to know how the trend of timing-shift goes

The cron runs the script on every Sunday 5:01AM and diagnostics will take about 5 min.

 

! Important:

During the running of the script, OMC and DARM can not be locked.

If you want to lock OMC and DARM in the early morning of weekend, just log in allegra and then comment out the command by using 'crontab -e'

 

 

  2152   Tue Oct 27 18:19:14 2009 robUpdateLockingbad

Quote:

Lock acquisition has gone bad tonight. 

The initial stage works fine, up through handing off control of CARM to MCL.  However, when increasing the AO path (analog gain), there are large DC shifts in the C1:IOO-MC_F signal.  Eventually this causes the pockels cell in the FSS loop to saturate, and lock is lost. 

 This problem has disappeared.  I don't know what it was. 

The first plot shows one of the symptoms.  The second plot is a similar section taken from a more normal acquisition sequence the night before.

All is not perfect, however, as now the handoff to RF CARM is not working.

Attachment 1: MCF_issue.png
MCF_issue.png
Attachment 2: no_MCF_issue.png
no_MCF_issue.png
  2151   Tue Oct 27 18:01:49 2009 robUpdatePSLhmmm

A 30-day trend of the PCDRIVE from the FSS.

Attachment 1: pcdrive_trend.png
pcdrive_trend.png
  2150   Tue Oct 27 17:58:25 2009 JenneConfigurationPEMUnknown PEM channels in the PEM-ADCU?

Does anyone know what the channels plugged in to the PEM ADCU, channels 5,6,7,8 are?  They aren't listed in the C1ADCU_PEM.ini file which tells the channel list/dataviewer/everything about all the rest of the signals which are plugged into that ADCU, so I'm not sure if they are used at all, or if they're holdovers from some previous time.  The cables are not labeled in a way that makes clear what they are.  Thanks!

  2149   Tue Oct 27 15:55:04 2009 KojiUpdateGeneralISS injection work / HEPA is on

I was working on the ISS excitation to take TFs.

I used ISS IL excitation, stealing from a small box on the floor for the OMC.

All the configuration was restored except that the HEPA is on.

  2148   Tue Oct 27 01:45:02 2009 robUpdateLockingMZ

Quote:
Tonight we also encountered a large peak in the frequency noise around 485 Hz. Changing the MZ lock point (the spot in the PZT range) solved this.


This again tonight.

It hindered the initial acquisition, and made the DD signal handoff fail repeatedly.
  2147   Mon Oct 26 23:14:08 2009 KojiUpdatePSLlaser power is down

I adjusted the steerings to the PMC and gained 7%. Now the MC_TRANS 7.0 has been recovered.

Actually I need another 7% to get MC_TRANS 7.5.
But I couldn't find how I can recover 126MOPA-AMPMON to 2.8ish.

Quote:

The laser power is down 5-6%

 

Attachment 1: PSL091026.png
PSL091026.png
  2146   Mon Oct 26 19:12:50 2009 kiwamuUpdateLSCdiagnostic script for LSC timing

The diagnostic script I've written is available in the 'caltech/users/kiwamu/work/20091026_OMC-LSC-diag/src'.

To run the script, you can just execute 'run_OMC_LSC.sh' or just call the m-file ' OMC_LSC_timinig.m'  from matlab.

 

NOTES:

The script destructs the lock of DARM and OMC, be careful if you are working with IFO.

  2145   Mon Oct 26 18:49:18 2009 kiwamuUpdateLSCOMC-LSC timing issue

According to my measurements I conclude that LSC-signal is retarded from OMC-signal with the constant retarded time of 92usec.
It means there are no timing jitter between them. Only a constant time-delay exists.

(Timing jitter)
Let's begin with basics.
If you measure the same signal at OMC-side and LSC-side, they can have some time delay between them. It can be described as followers.

exp1.png
where tau_0 is the time delay. If the tau_0 is not constant, it causes a noise of the timing jitter.

(method)
I have injected the sine-wave with 200.03Hz into the OMC-LSC_DRIVE_EXC. Then by using get_data, I measured the signal at 'OMC-LSC_DRIVE_OUT' and 'LSC-DARM_ERR' where the exciting signal comes out.
In the ideal case the two signals are completely identical.
In order to find the delay, I calculated the difference in these signals based on the method described by Waldman. The method uses the following expression.
exp2.png
Here the tau_s is the artificial delay, which can be adjusted in the off line data.
By shifting tau_s we can easily find the minimal point of the RMS, and at this point we can get tau_0=tau_s.
This is the principle of the method to measure the delay.  In my measurement I put T=1sec. and make the calculation every 1sec. in 1 min.


(results)
Attachment is the obtained results. The above shows the minimum RMS sampled every 1sec. and the below shows the delay in terms of number of shifts.
1 shift corresponds to Ts (=1/32kHz).  All of the data are matched with 3 times shift, and all of the minimum RMS are completely zero.
Therefore I can conclude that LSC-signal is retarded from OMC-signal with constant retarded times of 3*Ts exactly, and no timing jitter has been found.
 

Attachment 3: OMC_LSC60sec.png
OMC_LSC60sec.png
  2144   Mon Oct 26 18:15:57 2009 robUpdateIOOMC OLG

I measured the mode cleaner open loop gain.  It's around 60kHz with 29 degs of phase margin.

  2143   Mon Oct 26 17:45:34 2009 JenneUpdateAdaptive FilteringNew changes to the OAF fe code

[Alex, Jenne, Sanjit]

Alex came to the 40m today, and did several awesome things in OAF-land.

We discovered that there is, in fact, an ADC board connected to the ASS machine.  The tricky bit is that it only has a ribbon cable connector, so before we can use this ADC, we need to figure out how to make a breakout board/cable/something to connect the seismometer/accelerometer/microphone BNCs to this little board.  This is the same little board that connects the timing slave to the ASS machine.  For good or for ill, the timing slave is connected to this board via clip-doodles.  Potentially we can connect an ADC tester board to this board, and go from seismometer BNCs to clipdoodles to the tester board, but I'm not in love with the idea of utilizing clipdoodles as a semi-permanent solution until the upgrade.  I emailed Ben to see if he has a better idea, or (better yet) some spare hardware now that's the same as we'll use after the upgrade.  If we can use this ADC, it may solve our timing problem which is caused by the 110B ADC used by the PEM computer. Alex showed Sanjit and I how to connect the ASS's ADC card to the simulink diagram, when we're ready for that.

We also poked around in the code, and it seems that we can now save and restore OAF coefficients at will.  I added buttons to the OAF (ASS) screen, and Alex made it so the OAF coefficients are saved in RFM shared memory whenever you click the "save coeffs" button, and are restored when you click the "restore coeffs" button.  The buttons are the same as the 'Reset' button which has been there for a long time, so they seem to maybe have a similar problem in that you have to hold the button for a while in order for the code to realize that the button has been depressed.  We couldn't fix this easily, because it looks like our SimuLink cds stuff is a little out of date.  Some day (before/when Joe and Peter make new screens for the new 40m), we need to update these things.  Alex was concerned that it might take a while to do this, if the update broke some of the blocks that we're currently using.  Also, Sanjit and I now need to check that the coefficient-saving is going as planned.  When I have DTT open, and the OAF running, I see a certain shape to the signal which is sent to MC1 to correct for the seismic motion.  This shape includes at least several peaks at resonant frequencies that exist in our stacks/suspensions.  I can then save the coefficients, reset the active filter, and then restore the coefficients.  When I do this while watching DTT, it seems as though the general shape of the filter is restored, but none of the detailed features are.  The reason for this is still under investigation. 

The code-modifications involved a few iterations of 'remaking the ass'.

  2142   Mon Oct 26 15:40:01 2009 steveUpdatePSLlaser power is down

The laser power is down 5-6%

Attachment 1: laserpowerdown.jpg
laserpowerdown.jpg
  2141   Mon Oct 26 03:57:06 2009 robUpdateLockingbad

Lock acquisition has gone bad tonight. 

The initial stage works fine, up through handing off control of CARM to MCL.  However, when increasing the AO path (analog gain), there are large DC shifts in the C1:IOO-MC_F signal.  Eventually this causes the pockels cell in the FSS loop to saturate, and lock is lost. 

  2140   Sun Oct 25 14:29:45 2009 haixing, kiwamuConfigurationGeneralSR785 spectrum analyzer

In this morning, we have disconnected SR785 which was in front of 1X2 rack, to measure a Hall sensor noise.

After a while, we put back SR785 and re-connected as it has been.

But the display setup might have been changed a little bit.

 

  2139   Sat Oct 24 04:57:33 2009 ranaUpdateloremarconi phase

Quote:

So, it appears that one doesn't even have to change the Marconi set frequency to alter the phase of the output signal.  It appears that other front panel actions (turning external modulations on/off, changing the modulation type) can do it as well.  At least that's what I conclude from earlier this morning, when after setting up the f2 Marconi (166MHz) for external AM, the double-demod handoff in the DRMI no longer worked.  Luckily this isn't a real problem now that we have the setDDphases and senseDRM scripts. 

 The real problem is that we are using frequency synthesizers to make the beat signals (133 and 199) instead of mixers. Luckily, the future 40m will not use beat signals (?) or synthesizers.

  2138   Fri Oct 23 15:02:00 2009 robUpdateloremarconi phase

So, it appears that one doesn't even have to change the Marconi set frequency to alter the phase of the output signal.  It appears that other front panel actions (turning external modulations on/off, changing the modulation type) can do it as well.  At least that's what I conclude from earlier this morning, when after setting up the f2 Marconi (166MHz) for external AM, the double-demod handoff in the DRMI no longer worked.  Luckily this isn't a real problem now that we have the setDDphases and senseDRM scripts. 

  2137   Fri Oct 23 09:13:45 2009 steveSummaryVACRGA scan

Pump down #66 is 435 days old. RGA scan is normal. New maglev is fine. New UPS is in place but not hooked up to communicate.

V1 has bare minimum interlock. Pirani vacuum gauges  PTP1 and PRP do not communicate with readout system.

There is no emergency dial out in case of vacuum loss.  Our existing vacuum dedicated desk top computer is dead.

New cold cathodes, Pirani gauges and gauge controller should be added.

In general: vacuum system needs an upgrade !

 

Attachment 1: pd66md435.jpg
pd66md435.jpg
Attachment 2: pd66d435ptt.jpg
pd66d435ptt.jpg
  2136   Thu Oct 22 23:14:54 2009 ZachUpdateWIKI-40M UpdatePSL Table Diagram wiki entry

Quote:

Diagram. I don't want to say PNG is an editable format for this purpose...
You have the PPT, PDF or any drawing format to create this diagram.

Quote:

http://lhocds.ligo-wa.caltech.edu:8000/40m/PSL_Table_Diagram

Thanks. I love this. Could you also put the original file that is editable for future modification by anyone?

 Do you mean the diagram or the inventory? The diagrams are online as attachments (small versions on the main "PSL Table Diagram" page and large versions on the linked pages). The inventory is easily editable on the wiki itself. It's just rendered in table form using the CSV parse utility for "comma-separeted values" (though you actually need to use semicolons, for reasons unknown).

 

 

Good news and bad news. For the MOPA diagram, which I did recently, I have GIMP file with separate layers for the background image, ray traces, and labels. Unfortunately, I didn't realize that this was the best way to do it until I had done most of the ray tracing for the main diagram, so, although I have that file in GIMP as well, only the labels are on a separate layer. If this is a major issue I can do the tracing again. The other thing is that the original files are quite large: 17.3 MB for the MOPA, and 64.1(!) MB for the main diagram. Let me know what you think.

  2135   Thu Oct 22 21:58:26 2009 KojiUpdateWIKI-40M UpdatePSL Table Diagram wiki entry

Diagram. I don't want to say PNG is an editable format for this purpose...
You have the PPT, PDF or any drawing format to create this diagram.

Quote:

http://lhocds.ligo-wa.caltech.edu:8000/40m/PSL_Table_Diagram

Thanks. I love this. Could you also put the original file that is editable for future modification by anyone?

 Do you mean the diagram or the inventory? The diagrams are online as attachments (small versions on the main "PSL Table Diagram" page and large versions on the linked pages). The inventory is easily editable on the wiki itself. It's just rendered in table form using the CSV parse utility for "comma-separeted values" (though you actually need to use semicolons, for reasons unknown).

 

  2134   Thu Oct 22 15:49:29 2009 ZachUpdateWIKI-40M UpdatePSL Table Diagram wiki entry

Quote:

http://lhocds.ligo-wa.caltech.edu:8000/40m/PSL_Table_Diagram

Thanks. I love this. Could you also put the original file that is editable for future modification by anyone?

Quote:

 I made a wiki entry for the PSL table diagram under the PSL directory on the 40mHomePage. I tried to use the ImageLink macro to use a resized (smaller) version of the diagram as a link to the full image, which it is designed to do if there is no target given, but it didn't seem to work. Instead, I had to create a second page that had the full-sized diagram, and I used ImageLink with a smaller version to link to that page.

The inventory that is shown is clearly incomplete. Part of this is due to the fact that many labels were either missing or impossible to read without touching stuff. For those components with labels missing, I tried to infer what they were to the best of my knowledge, but I wasn't able to for all of them. In true wiki spirit, everyone is encouraged to fill in any additional information they might have on these components. 

 

 Do you mean the diagram or the inventory? The diagrams are online as attachments (small versions on the main "PSL Table Diagram" page and large versions on the linked pages). The inventory is easily editable on the wiki itself. It's just rendered in table form using the CSV parse utility for "comma-separeted values" (though you actually need to use semicolons, for reasons unknown).

  2133   Thu Oct 22 15:44:16 2009 ZachUpdateWIKI-40M UpdateMOPA diagram

 I have updated the PSL Diagram wiki page to include MOPA. As with the PSL diagram, clicking the photo on the main page takes you to a larger image. The inventory is pretty meager as I didn't have time to sit and read labels (if indeed there are any). I will look through the documentation at the 40m to see if there is a record of what is there. Again, if you know something, please amend the list!!

http://lhocds.ligo-wa.caltech.edu:8000/40m/PSL_Table_Diagram

  2132   Thu Oct 22 08:45:58 2009 steveUpdateIOOIP ang & pos recentered

Quote:

Pointing stability of 4 days. Initial pointing does not go through suspended optics. It is launched  right after the Piezo Jena steering mirrors in the BS chamber.

IP-ANG on epics screen is  C1:ASC-IBQPD_X and Y in dataviewer  were recentered. This beam is clipping a bit in ETMX chamber  pick off mirror.

IP-POS pick  off is in the BS chamber and it's qpd on the BS_ISCT This beam is also clipping just a little bit. This is easy to fix. We'll have to remove an iris from the BS optical levers table.

note: arms were not locked when I recentered

 IP-ANG clipping can be traced back to our last vent of Aug. 18, 2008  See elog entry #845

This was an after earth quake - sus repair vent

Attachment 1: pointing1000.jpg
pointing1000.jpg
  2131   Wed Oct 21 17:12:30 2009 AlbertoUpdateelogBrowser context menu enabled on the Elog under HTM editing mode

On behalf of Steve and of the rest of the not-native-English community at the 40m willing to have their browser's spell checker working while editing the Elog, I fixed the Elog's feature that prevented Firefox' context menu (that one which pops up with a mouse right click) to work when using the HTML editing interface (FCKeditor).

That let also Firefox spell checker to get enabled.

To get the browser context menu just press CTRL right-clicking.

To make sure that the features works properly on your browser, you might have to fully clear the browser's cache.

Basically I modified the FCKeditor config file (/cvs/cds/caltech/elog/elog-2.7.5/scripts/fckeditor/fckconfig.js). I added this also to the elog section on our Wiki.

ELOG V3.1.3-